The power, P, produced by a wind turbine is given by:

P = ( ½) C𝜌Av3

where 𝜌 is the density of air [ML-3 ], A is the cross-sectional area [L2 ] of the blade and v is the velocity of air. Determine in MLT the dimensions of:

i. v3

ii. P

iii. C


1
Expert's answer
2021-06-11T16:58:50-0400

i)

"v^3=(LT^{-1})^3=L^3T^{-3}"


II)

"P=ML^2T^{-3}"


III)

"C=\frac{ML^2T^{-3}}{ML^{-3}\cdot L^3T^{-3}L^2}=1"